HOP# 1 Job #AAI 10, which was completed during January, cosas Job #AAI 11. which was also completed during January, consisted of 70 units. Jobs #AAI 12 and #AAI 13 were not coupleted during January The Job Shop sells units for 140% of their manufacturing cost. January, cousisted of 40 uuts. REQUIRED a. Prepare journal eatries to record the following:w Ca 1. Use of direct materials during Janry 2. The direct labor payroll for the month. 3. Application of mamfacturing ovethead for J 4. Actual manufacturing overhead for January 5. Completion of Jobs #AAI 10 and #AA111 6. A credit sale of 30 units from Job #AAI 10 and 40 January units from Job #AAI 11. b. What is The Job Shops over- or under-applied overhead at the end of January? c. What was the balance of The Job Shops Work-in-Process account on Jaumary 1 20 8? d. What was the balance of The Job Shops Work-in-Process account on January 31 20 8? )lo

HOP#1 The Job Shop prodtuces a variety of products nsing the equipment in their factory. They use job order costing and apply manufacturing overhead on the basis of machine hours worked using a predetertinined overfiead rate. During 20 S, The Job Shop expects to incur S260,000 of minfactnring overhead. They expect to work 10,400 machine hours On January 1, 20-8. The Job Shop had one Job #AA 110 in the factory. The job cost sheet for Job #AAI 10 shows that $5,400 of direct material had been used on the Job 20 7. Direct labor used on the job in 20 7 was S1.050 and manufacturing over applied to the job in 20_7 was $3,450. Production was started on threejobs during January: #AA111, #AAI 12. #AA113. Material requisitioned during January was as follows $800 for #AA110 $6,700 for #AA111 S4.900 for #AAI 12 S3,100 for HAA13 Direct labor incutred during January was as follows S300 for #AA110 $1.500 for #AA111 $900 for #AA112 S500 for #AA113 Machine hours worked during January were as follows 80111h on #AA110 260th on #AA111 150mil ou mAAI 12 I 101nh on #AA113 Manufacturing Overhead incurred during Jannary was as follows: $13.600 of depreciation on plaat and equipment S500 of insurance on the plant $3.100 of utilities, and $3,700 of indirect labor. Job Shop had paid for the insurance in 20 7. Utilities will be paid for in February 20 8.
Answer #1
a. Sl No. Account titles and explanation Debit Credit
1 Work in process control account 15500
Material control account 15500
(800+6700+4900+3100)
(Direct materials used)
2 Work in process control account 3200
Wages control account 3200
(300+1500+900+500)
(Direct labor payroll for the month)
3 Work in process control account (Note:1) 15000
Manufacturing overhead control account 15000
(Manufacturing overhead applied)
4 Manufacturing overhead control account 20900
Accumulated depreciation 13600
Prepaid insurance (Since paid in Previous year) 500
Utilities payable (Since to be paid in Feb) 3100
Wages control account 3700
(Manufacturing overhead incurred)
5 Finished goods control account (Note:2) 17800
Work in process control account 17800
(Completion of Jobs)
6 Accounts receivable (Note:3) 15015
Sales 15015
(Sales on credit)
Cost of goods sold (Note:4) 10725
Finished goods control account 10725
(Cost of sales recorded)
Note:1
Predetermine overhead rate=Estimated manufacturing overhead/Estimated machine hours=260000/10400=$25 per machine hour
Overhead applied=Machine hours worked* Predetermined overhead rate
Machine hours worked:
mh
#AA110 80
#AA111 260
#AA112 150
#AA113 110
Total 600
Overhead applied=600* 25=$ 15000
Note:2
Cost of completed Jobs:
#AA110 #AA111 Total
Direct material 800 6700
Direct labor 300 1500
Manufacturing overhead 2000 6500
(Machine hours worked* Predetermined overhead rate) (80*25) (260*25)
Total 3100 14700 17800
Note:3
#AA110 #AA111 Total
Cost of jobs 3100 14700
Add: 40% markup 1240 5880
(140% of mfg cost means 40% of markup)
Selling price a 4340 20580
Units produced b 40 70
Selling price per unit c=(a/b) 108.5 294
Units sold d 30 40
Sales value e=(c*d) 3255 11760 15015
Note:4
#AA110 #AA111 Total
Cost of jobs a 3100 14700
Units produced b 40 70
Cost per unit c=(a/b) 77.5 210
Units sold d 30 40
Cost of sales e=(c*d) 2325 8400 10725